if a lenth of simple pendulam is increse by 2%, the the time period

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Abhishek
 
T = 2π√(L/g)

From above
ΔT/T × 100 = 0.5 × (ΔL/L × 100)
= 0.5 × 2%
= 1%
∴ If length of pendulum is increased by 2% then time period will increase by 1%
Hope it helps
 
Regards
Arun